# Conformal prediction for causal effects of continuous treatments

### v2

@@ E.1 Synthetic dataset generation
E.2 Medical dataset
We use the MIMIC-III dataset (Johnson et al., 2016), which includes electronic health records (EHRs) from patients admitted to intensive care units. From this dataset, we extract 8 confounders (heart rate, sodium, blood pressure, glucose, hematocrit, respiratory rate, age, gender) and a continuous treatment (mechanical ventilation) using an open-source preprocessing pipeline (Wang et al., 2020). From each patient trajectory in the EHRs, we sample random time points and average the value of each variable over the ten hours before the sampled time point. We define the variable blood pressure after treatment as the outcome, for which we additionally apply a transformation to be more dependent on the treatment and less on the blood pressure before treatment. We remove all patients (samples) with missing values and outliers from the dataset. Outliers are defined as samples with values smaller than the 0.1th percentile or larger than the 99.9th percentile of the corresponding variable. The final dataset contains 14719 samples, which we split into train (60%), val (10%), calibration (20%), and test (10%) sets.
E.3 Implementation details
Our experiments are implemented in PyTorch Lightning. We provide our code in our GitHub repository.
@@ E.3 Implementation details
Appendix F Further results
@@
Performance: We first report the performance of the underlying prediction models $\phi$ for the synthetic datasets across 10 runs in Table 2. The prediction model on the real-world dataset achieved a mean squared error loss of 1.2373.
@@
Table 2: Mean and standard deviation of MSE loss of prediction models $\phi$ across 10 runs.

### v3

@@ F.1 Synthetic dataset generation
F.2 Medical dataset
We use the MIMIC-III dataset [26], which includes electronic health records (EHRs) from patients admitted to intensive care units. From this dataset, we extract 8 confounders (heart rate, sodium, blood pressure, glucose, hematocrit, respiratory rate, age, gender) and a continuous treatment (mechanical ventilation) using an open-source preprocessing pipeline [55]. From each patient trajectory in the EHRs, we sample random time points and average the value of each variable over the ten hours before the sampled time point. We define the variable blood pressure after treatment as the outcome, for which we additionally apply a transformation to be more dependent on the treatment and less on the blood pressure before treatment. We remove all patients (samples) with missing values and outliers from the dataset. Outliers are defined as samples with values smaller than the 0.1th percentile or larger than the 99.9th percentile of the corresponding variable. The final dataset contains 14719 samples, which we split into train (60%), val (10%), calibration (20%), and test (10%) sets.
F.3 Implementation details
Our experiments are implemented in PyTorch Lightning. We provide our code in our GitHub repository. All experiments were run on an AMD Ryzen 7 PRO 6850U 2.70 GHz CPU with eight cores and 32GB RAM.
@@ F.3 Implementation details
F.4 Selection of the interventions in our experiments
The treatment (in the complete dataset) is modeled to lie in the range [0,40]. Therefore, the treatments/interventions can also only fall into this range. All samples that would have achieved a treatment outside this range through the interventions were neglected in our analysis. To guarantee that a sufficient number of samples were included in our experiments, we chose the maximal soft treatment as an increase of 10.
[…]
Appendix G Further results
@@
Performance: We first report the performance of the underlying prediction models $\phi$ for the synthetic datasets across 50 runs in Table 4. The prediction model on the real-world dataset achieved a mean squared error loss of 1.2373.
@@
Table 4: Mean and standard deviation of MSE loss of prediction models $\phi$ across 50 runs.
We further report the width of the prediction intervals in our synthetic experiments in Table 5. The width is important to assess the usefulness of the resulting prediction intervals. As the performance of the ensemble method is not comparable with the coverage of MC-Dropout and our CP method, we only compare the latter two methods with regard to the interval width.
Dataset 1 Dataset 2
Delta Ours MC-Dropout Ours MC-Dropout
1 0.3647 (0.1284) 0.1938 (0.1170) 0.4051 (0.1036) 0.2897 (0.1480)
5 0.4024 (0.2285) 0.1653 (0.1103) 0.4610 (0.2479) 0.3036 (0.1455)
10 0.4301 (0.2610) 0.1639 (0.1080) 0.6711 (0.8520) 0.3235 ( 0.1445)
Table 5: Mean and standard deviation of the resulting prediction intervals.